A wave has a length of 2m and a frequency of 10Hz. Calculate the wave number . Calculate the angular frequency. Calculate wave speed


1
Expert's answer
2021-07-07T12:52:21-0400

Gives

n=10Hz

"w=\frac{2\pi}{T}"

"T=\frac{1}{n}=\frac{1}{10}"

T=0.1Hz

"w=\frac{2\pi}{T}=\frac{2\pi}{0.1}=20π rad/sec"


Wave number

"k=\frac{2\pi}{\lambda}"

"k=\frac{2\pi}{2}=π" "m^{-1}"

Speed wave

"v=\frac{w}{k}=\frac{20π}{π}=20m/sec"